Axolotl:

I need quick help with this question, please! It shouldn't be to hard. Ray IL bisects angle HIJ. If m∠HIL = (6x − 7)° and m∠JIL = (5x + 4)°, what is m∠JIL?

Axolotl:

I got 11, but I have a feeling I should be doing more....

Axolotl:

@jhonyy9

SwaggyMark:

11 correct

Axolotl:

@swaggymark wrote:
11 correct
ok thank you

SwaggyMark:

yea because the bisection make em congruent and u put the equation equal to each other and get 5x + 4 = 6x - 7 , then u minus 4 and get 5x = -11 + 6x , then you gotta minus 6x and get -x = -11 and once you fraction the -11 and -1 you get x = 11